# Methodology for Fine-Grain GPU Power Visibility and Insights

### v1

@@ VIII Conclusion
We focus in this work on fine-grain profiling for key AI operations of matrix-matrix multiplication and communication kernels on GPUs, which are widely deployed for AI workloads. To this end, we first identified challenges in doing fine-grain GPU power profiling and proposed FinGraV methodology to address these challenges on a state-of-the-art AMD MI300X GPU. We identify several important takeaways from FinGraV profiles and conclude by making many recommendations to optimize GPU power including algorithmic/software techniques that can lower power, the necessity of focusing on lowering power for compute-light GPU kernels, and the necessity for smarts in power management infrastructure.
@@ Acknowledgment
AMD, the AMD Arrow logo, AMD CDNA, AMD Instinct, AMD ROCm, AMD Infinity Cache, AMD Infinity Fabric, and combinations thereof are trademarks of Advanced Micro Devices, Inc. Other product names used in this publication are for identification purposes only and may be trademarks of their respective companies.

### v2

@@ VIII Conclusion
We focus in this work on fine-grain profiling for key AI operations of matrix-matrix multiplication and communication kernels on GPUs, which are widely deployed for AI workloads. To this end, we first identified challenges in doing fine-grain GPU power profiling and proposed FinGraV methodology to address these challenges on a state-of-the-art AMD MI300X GPU. Our work identifies important power measurement guidance to avoid steep measurement errors (as steep as 80%) in power (and hence energy). Additionally, we identify several important takeaways from FinGraV power profiles which in turn lead to power optimization recommendations focused on GPU sub-component power consumption and GPU power proportionality.
@@ Acknowledgment
The authors thank Nuwan Jayasena, Ashish Jain, Steve Kushnir, Aranyak Mishra, Yuri Lee, and the anonymous ISPASS reviewers for helping improve the paper. AMD, the AMD Arrow logo, AMD CDNA, AMD Instinct, AMD ROCm, AMD Infinity Cache, AMD Infinity Fabric, and combinations thereof are trademarks of Advanced Micro Devices, Inc. Other product names used in this publication are for identification purposes only and may be trademarks of their respective companies.